OpenStudy (anonymous):

Solve the equation: x2 - 12x = -36 explain your answer

OpenStudy (diyadiya):

\[x^2-12x=-36 \]Add 36 on both sides what do you get?

OpenStudy (anonymous):

in order to solve these equations they need to be equal to 0, so first you need to add 36 to both sides

OpenStudy (maheshmeghwal9):

\[x^2-12x+36=0\]

OpenStudy (maheshmeghwal9):

\[x^2-6x-6x+36=0\]

OpenStudy (maheshmeghwal9):

\[(x-6)(x-6)=(x-6)^2\]=0

OpenStudy (maheshmeghwal9):

hence ,x=6.